OpenStudy (anonymous):

which of the following is a polynomail? a. x^-x+2 b.x^-2-1 c.x^4-2 d.(x^6-4)(x^-5+10

OpenStudy (solomonzelman):

a polynomial is a function where variable(s) is/are raised to the power, and this power is a whole number.

OpenStudy (solomonzelman):

Some examples of a polynomial is: \(\\[1.2em]\) \(\large\color{black}{ \displaystyle y=0 \\[0.7em] }\) \(\large\color{black}{ \displaystyle y=-34 \\[0.7em] }\) \(\large\color{black}{ \displaystyle y=x^2+3 \\[0.7em] }\) \(\large\color{black}{ \displaystyle y=-x-4 \\[0.7em] }\) \(\large\color{black}{ \displaystyle y=12x^6-7x^5+2x^4+3x^3-x^2-x+3 \\[0.7em] }\)